Using the following chemical equation how many grams of carbon dioxide can be produced by burning 18.49 g of butane?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-18T07:29:32-0500

According to the equation, equal numbers of moles of butane and carbon dioxide must be used and produced, respectively. From here:

m(CO2) = m(C4H10) × Mr(CO2) / Mr(C4H10) = 18.49 g × 44.01 g / 58.12 g = 12.76 g


Answer: 12.76 g